OpenStudy (anonymous):

a line passes through the point (1,2) and has a slope of 9. write an equation for this line

OpenStudy (yttrium):

You can use point slope form. \[y-y _{1} = m(x-x _{1})\] Where m = slope and \[x _{1} and y _{1}\] are the coordinates of the given point.

OpenStudy (anonymous):

ok so far i have y-2=9x-9

OpenStudy (yttrium):

Yes. But that is not the final answer yet. You have to convert it to standard form. You know it?

OpenStudy (anonymous):

y=9x-7

OpenStudy (yttrium):

Got it. :)